OpenStudy (anonymous):

solve with a quadratic equation. the product of two consecutive odd integers is 2, 115. Find the integers.

OpenStudy (anonymous):

x + x+1= 2,115, solve for x ^_^ consecutive integers are : x, x+1, x+2, x+3 etc

OpenStudy (anonymous):

wait, is it x? or x+1 as a starting point lol @_@

OpenStudy (amistre64):

ax^2 + bx + c = 2 a(x+2)^2 + b(x+2) + c = 115

OpenStudy (amistre64):

what do the 2 and 115 mean? that (x)(x+2) = 2 and 115?

OpenStudy (anonymous):

im not sure, it just puts 2, 115 :-/

OpenStudy (amistre64):

or the (2+n)(2+(n+2)) = 115?

OpenStudy (amistre64):

a(2^2) + b(2) + c = 115 is what I think it means

OpenStudy (amistre64):

when x=2; y=115

OpenStudy (amistre64):

just cant seem to make sense of the question.. :)

OpenStudy (anonymous):

ah, well, maybe I'll just wait until monday and ask my teacher about it :P thanks for trying to help tho!

OpenStudy (amistre64):

youre welcome :) wish I coulda helped out more tho ;)

OpenStudy (anonymous):

i guess it must be product of two odd integers is 115

OpenStudy (amistre64):

just for kicks, let me try this... (2+n)(2+(n+2)) = 115 (2+n)(4+n) = 115 n^2 +6n +8 = 115 n^2 +6n -107 = 0 n= -6/2 +- sqrt(36 + 428)/2 n = -3 +- sqrt(464)/2 ; but that doesnt pruduce 2 consecutive odd integers for me...

OpenStudy (amistre64):

n(n+2) = 115 n^2 +2n = 115 n^2 +2n -115 = 0 n= -4/2 +- sqrt(4 + 230)/2 n = -2 +- sqrt(234)/2...which isnt 2/2 so that busts as well.....

OpenStudy (anonymous):

yea i tried ...same result

OpenStudy (anonymous):

there is something missing

OpenStudy (amistre64):

2,115 seems like a pair of coordinates.... but that doesnt make sense to me either...

OpenStudy (amistre64):

does f(2) = 115? and f(-2)=115? but then where doe we get the two consecutive number to play into it?

OpenStudy (amistre64):

if anything the 2 roots would multiply to get a number; but it aint 115..and it aint 2 that I can see

OpenStudy (anonymous):

did u try......product=2115???

OpenStudy (amistre64):

no, but then again I didnt try product = 47 either :)

OpenStudy (anonymous):

y 47?

OpenStudy (amistre64):

lol..... why 2115?

OpenStudy (anonymous):

like some ppl write 2115 as 2,115 :)

OpenStudy (anonymous):

i wasnt making a fun of u :)

OpenStudy (amistre64):

ahhhh!! damn yankees lol

OpenStudy (anonymous):

(n-47)(n+45)=0

OpenStudy (anonymous):

awwwwwww

OpenStudy (anonymous):

wrong :(

OpenStudy (amistre64):

(n)(n+2) = 2115 n^2 +2n -2115 = 0 n = -4/2 +- sqrt(4 + 8460)/2 n= -2 +- sqrt(8464)/2....

OpenStudy (anonymous):

hmmmm

OpenStudy (amistre64):

n = -2 +- 92/2

OpenStudy (amistre64):

its either 2 negs or 2 pos to get 2115 right?

OpenStudy (anonymous):

n^2+2n-2115=0 (n+47)(n-45)=0

OpenStudy (anonymous):

n=47 47+2=49

OpenStudy (anonymous):

yes right

OpenStudy (amistre64):

47(49)=2303

OpenStudy (anonymous):

yes :(

OpenStudy (amistre64):

its just a broken question ;)

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

but where is the real prob??when the factorization is correct

OpenStudy (anonymous):

oh 45 and 47 is the right choice

OpenStudy (anonymous):

likewise -45 and -47

OpenStudy (anonymous):

so product =2115 is the actual problem :)

OpenStudy (amistre64):

45 and 47 produce 2115; yes :)

OpenStudy (anonymous):

m so bad in calculations :P

OpenStudy (anonymous):

so it was just a wrong (,) there :P

OpenStudy (amistre64):

you did good :)

OpenStudy (anonymous):

so starlit u dont need to wait for monday...the solution is there for u

OpenStudy (anonymous):

nooooo, factoriztion was correct,but.........